# Very High-order Compact Gas-kinetic Scheme With Discontinuity Feedback Factor

### v1

@@ 4.10 High Mach number astrophysical jet
As shown in Fig. 15 and Fig. 16, the ASE-DFF(5,3)-CGKS, ASE-DFF(7,5,3)-CGKS and ASE-DFF(9,7,5,3)-CGKS have identified similar large-scale characteristic structures. Generally speaking, this type of problem has very high requirements for the positivity of the flux solver and the robustness of the reconstruction method. The conventional high-order GKS or CGKS cannot pass this problem. However, the robustness of the new CGKS method based on the discontinuity feedback factor has been greatly improved. At the same time, the resolution of the higher-order CGKS has also been relatively improved.
Figure 15: High Mach number astrophysical jet: density (left) and pressure (right) contours from ASE-DFF(5,3)-CGKS (top), ASE-DFF(7,5,3)-CGKS (middle) and ASE-DFF(9,7,5,3)-CGKS (bottom) in logarithmic scale. The Mach number is 80 and the resolution is $448\times 224$.
@@ 5 Conclusion
This paper presents a novel very high-order compact gas-kinetic scheme with adaptive stencil extension using discontinuity feedback factor. The scheme retains two-stage fourth-order time-stepping method of the original compact gas-kinetic scheme and updates the conservative variables and their gradients in the cell in each stage. The CGKS with the discontinuity feedback factor can achieve fifth-order, seventh-order and ninth-order accuracy. For compact stencils, the fifth-order compact stencil is same to a third-order non-compact stencil, while the seventh-order and ninth-order compact stencils are same to a fifth-order non-compact stencil. The elimination of the need to compute smoothness indicators for high-order reconstruction polynomials reduces computational costs. Additionally, a new two-dimensional reconstruction framework simplifies the construction of the non-equilibrium and equilibrium state required for CGKS flux evolution. This method aligns with commonly used high-order gas-kinetic scheme and constructs all equilibrium state using a kinetic weighting method. One-dimensional and two-dimensional accuracy tests demonstrate that the CGKS based on the discontinuity feedback factor can achieve up to ninth-order accuracy. Furthermore, a series of one-dimensional and two-dimensional test cases involving shock and rarefaction waves show that the compact high-order gas-kinetic scheme based on the discontinuity feedback factor possesses excellent shock-capturing capabilities and robustness. Its low dissipation characteristics enable high resolution at discontinuities. The CFL time step used in the numerical tests remains relatively large. The new and simplified CGKS based on the discontinuity feedback factor achieves up to ninth-order accuracy, demonstrating significant potential for achieving higher accuracy on relatively compact stencils in future unstructured mesh applications.
@@ Acknowledgment
The first author is very grateful to Haolin Liu and Qingdian Zhang for all the help provided to me. And the authors would like to thank Yixiao Wang, Hongyu Liu, Cong shi, Qihui Gao, YaLing Wu and Zhiwen Zhuang for helpful discussion. This work was supported by the National Natural Science Foundation of China (Nos. 12302378, 92371201, and 12172316), and the Funding of National Key Laboratory of Computational Physics, National Key R&D Program of China (Grant Nos. 2022YFA1004500)

### v2

@@ 4.10 High Mach number astrophysical jet
As shown in Fig. 15 and Fig. 16, the ASE-DFF(5,3)-CGKS, ASE-DFF(7,5,3)-CGKS and ASE-DFF(9,7,5,3)-CGKS have identified similar large-scale characteristic structures. Generally speaking, this type of problem has very high requirements for the positivity of the flux solver and the robustness of the reconstruction method. For non-compact GKS, this problem cannot be solved no matter what discontinuity feedback factor is used [48]. For compact-GKS, even without using the discontinuity feedback factor, this test cannot be passed. Therefore, the robustness of the new CGKS method based on the discontinuity feedback factor has been greatly improved. At the same time, the resolution of the higher-order CGKS has also been relatively improved. Furthermore in Fig. 15, compared with ASE-DFF-CGKS, the non-compact results based on ASE-DFF-LF not only have the lower resolution at the top of the jet, but also exhibit greater oscillation in the flow field. In Fig. 16,the density and pressure contours of high Mach number astrophysical jet from ASE-DFF-CGKS is shown and remain satisfactory.
Figure 15: High Mach number astrophysical jet: density contours from non compact scheme with LF(left) and compact gas-kinetic scheme(right): ASE-DFF(5,3) (top), ASE-DFF(7,5,3) (middle) and ASE-DFF(9,7,5,3) (bottom) in logarithmic scale. The Mach number is 80 and the resolution is $448\times 224$.
@@
[…]
Table 7: The CPU time(s) of reconstruction process for the ASE-DFF-CGKS and the HWENO-AO-CGKS.
Mesh size ASE-DFF(5,3)-CGKS ASE-DFF(7,5,3)-CGKS ASE-DFF(9,7,5,3)-CGKS
$\times100$ 0.913 1.382 1.942
$\times200$ 5.244 6.048 7.685
$\times400$ 15.764 28.722 30.228
$\times800$ 63.766 126.264 119.484
Mesh size HWENO(5,3)-CGKS HWENO-AO(7,5,3)-CGKS HWENO-AO(9,5,3)-CGKS
$\times100$ 1.187 3.105 4.448
$\times200$ 6.589 15.064 17.069
$\times400$ 24.146 63.387 67.848
$\times800$ 97.128 256.688 284.287
The discontinuity feedback factor not only has the less computational cost but also enhances the ability of schemes to solve shock waves and rarefaction waves. As shown in Table  8,, the test results of HWENO-AO-CGKS and ASE-DFF-CGKS are summarized. It can be seen that in the Le Blanc problem and the high Mach number astrophysical jets with Ma = 80 and 2000, ASE-DFF-CGKS can pass the test. But HWENO-AO-CGKS fails to pass these two extreme cases.
Table 8: Robustness verification summary of both HWENO-type and DFF-enhanced CGKSs.
Test case Scheme HWENO-AO-CGKS ASE-DFF-CGKS
Shock-tube problem $\checkmark$ $\checkmark$
Shock-density wave interaction $\checkmark$ $\checkmark$
Interacting blast waves $\checkmark$ $\checkmark$
Double rarefaction wave problem $\checkmark$ $\checkmark$
Le Blanc problem $\times$ $\checkmark$
Two-dimensional Riemann problems $\checkmark$ $\checkmark$
Double Mach reflection problem $\checkmark$ $\checkmark$
Viscous shock tubes problem $\checkmark$ $\checkmark$
High Mach number astrophysical jet $\times$ $\checkmark$
@@ 5 Conclusion
This paper presents a novel, very high-order compact gas-kinetic scheme (CGKS) featuring adaptive stencil extension with a discontinuity feedback factor. The proposed scheme retains the two-stage fourth-order time-stepping method of the original CGKS and updates both the conservative variables and their gradients within each cell at every stage. With the incorporation of the discontinuity feedback factor, the CGKS achieves fifth-, seventh-, and ninth-order spatial accuracy. Notably, for compact stencils, the fifth-order compact stencil is equivalent in size to a third-order non-compact stencil, while seventh- and ninth-order compact stencils correspond to those of a fifth-order non-compact scheme. The elimination of the need to compute smoothness indicators for high-order reconstruction polynomials significantly reduces computational cost. Furthermore, a new two-dimensional reconstruction framework simplifies the construction of non-equilibrium and equilibrium states required for the CGKS flux evolution. This approach is consistent with commonly used high-order gas-kinetic schemes, utilizing a kinetic weighting method to construct all equilibrium states. One- and two-dimensional accuracy tests demonstrate that the CGKS with the discontinuity feedback factor can achieve up to ninth-order accuracy. Moreover, a series of test cases involving shock and rarefaction waves confirm that this compact, high-order gas-kinetic scheme possesses excellent shock-capturing capabilities and robustness. Its low-dissipation characteristics enable high resolution at discontinuities, and the CFL time step used in numerical tests remains relatively large. Overall, the newly developed and simplified CGKS based on the discontinuity feedback factor achieves up to ninth-order accuracy, demonstrating significant potential for attaining even higher accuracy on compact stencils in future applications involving unstructured meshes.
@@ Acknowledgment
The authors would like to thank Haolin Liu, Qingdian Zhang, Yixiao Wang, Hongyu Liu, Cong shi, Qihui Gao, YaLing Wu and Zhiwen Zhuang for helpful discussions. The current research is supported by the National Natural Science Foundation of China (12302378, 12172316, 92371201, and 92371107), the Funding of National Key Laboratory of Computational Physics, the National Key R&D Program of China (Grant No. 2022YFA1004500), the Natural Science Basic Research Plan in Shaanxi Province of China (No. 2025SYS-SYSZD-070), and the Hong Kong Research Grant Council (16301222, and 16208324).